sum of values of x and y satisfying the equation 3– 4y = 77; 3x/2 – 2y = 7 is:

Grade:12

1 Answers

Arun
5 years ago
Dear shubham
 
let 3^(x/2) = t and 2^y = r
 
then
 
t – r = 7
 
t^2 – r^2 = 77
 
t + r = 11
 
hence
 
t-r = 7
t + r = 11
 
2t = 18
 
t = 9
and 
r = 2
 
hence
 
3^(x/2) = 9
 
x = 4
 
y = 1